Horsepower Hookup, Inc., is a large automobile company that specializes in the production of high-powered trucks. The company is determining cost allocations for purposes of performance evaluation. A portion of company bonuses depends on divisions achieving cost management goals. This necessitates highly accurate support department cost allocation. Management has also stated that it has the means to implement as complex a method as necessary.

The general manager over the Mid-Size D wants to get a good idea of what factors are driving the costs of the support departments in order to make accurate cost allocations, so finding accurate support department cost drivers is important. Support department costs include Janitorial ($163,100) and Security ($285,400). The Janitorial costs vary depending on the number of vehicles produced, increasing with larger production volumes. Security costs are fixed based on the size of the lot, and do not change with respect to how many vehicles are in the lot or warehouse. Joint costs involved in producing the trucks before the split-off point where the various makes, models, and colors are produced are $946,000 for the period. All makes, models, and colors sell at relatively similar margins, but the sports models and metallic colors are normally more difficult to produce during the joint production process.

a. Identify the support department cost allocation method that matches the given situation.

a. Direct method.

b. Reciprocal services method.

c. Sequential method.

d. None of the above.

b

b. Identify the cost driver that should be used to allocate janitorial costs.

a. square feet.

b. number of vehicles produced.

c. number of employees.

d. none of the above.

b

c. Identify the cost driver that should be used to allocate security costs.

a. number of vehicles produced.

b. number of employees.

c. size of the lot.

d. none of the above.

c

d. If Janitorial costs were to be allocated based on square footage, and Security costs based on asset value, what percentage of each support department’s costs would be allocated to each production department using the sequential method (allocating Security costs first) given the following:


Square
Footage
Asset
Value
Janitorial Department 3,000 $ 10,000
Security Department 2,000 2,300
Production Department 1 54,000 450,000
Production Department 2 36,000 540,000
Security Department cost allocation:
Janitorial Department: %
Production Department 1: %
Production Department 2: %
Janitorial Department cost allocation:
Production Department 1: %
Production Department 2: %
Answer #1

A. According to the description given above , the most accurate but complex method is required hence it will be the reciprocal method. Answer is option B.

B. Answer is option B. Number of vehicle produced.

C. Answer is option C. Size of the lot.

D. Security department cost allocation

Janitorial department = $10000 / ($10000+$450000+540000)

= 1%

Production department 1 = $450000/$1000000 = 45%

Production department 2 = $540000/$1000000 = 59%

Janitorial department cost allocation

Production department 1 = 54000/(54000+36000) = 60%

Production department 2 = 36000/90000 = 40%